# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v3

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message-passing) layers, within which the representation of each node is updated based on those of its neighbors. The most expressive message-passing GNNs can be obtained through the use of the sum aggregator and of MLPs for feature transformation, thanks to their universal approximation capabilities. However, the limitations of MLPs recently motivated the introduction of another family of universal approximators, called Kolmogorov-Arnold Networks (KANs) which rely on a different representation theorem. In this work, we compare the performance of KANs against that of MLPs on graph learning tasks. We evaluate two different implementations of KANs using two distinct base families of functions, namely B-splines and radial basis functions. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our results indicate that KANs are on-par with or better than MLPs on all studied tasks, making them viable alternatives, at the cost of some computational complexity. All code will be made available upon acceptance.
@@ 1 Introduction
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [1]. Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [2]. Most GNNs consist of a series of message-passing layers. Within a message-passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
@@
A lot of recent work has focused on investigating the expressive power of GNNs [3]. There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [4]. Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [5]. This line of work gave insights into the limitations of different models [6, 7], but also led to the development of more powerful architectures [4, 8, 9].
@@
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [10, 11]. The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Recently, Kolmogorov-Arnold Networks (KANs) [12] have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem [13] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs with learnable activations and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions and in settings where regularity is expected.
@@
In this paper, we present a thorough empirical comparison of the performance between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our intuition is that KANs could exploit the expected regularity underlying the message-passing paradigm to provide comparatively good performance on graph tasks. Our work is orthogonal to prior work that studies the expressive power of GNNs since we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the models’ performance. We evaluate the different GNN models on several standard node classification, graph classification, and graph regression datasets.
@@
The rest of this paper is organized as follows. Section 2, provides an overview of the tasks we address in this paper, as well as a description of message-passing GNNs and Kolmogorov-Arnold networks. In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In Section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks. Finally, Section 7 concludes the paper.
@@ 2.1 Graph Learning Tasks
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn how to map nodes to their class labels, i. e., to learn a function $f_{\texttt{node}}$ such that $f_{\texttt{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., a function $f_{\texttt{graph}}$ such that $f_{\texttt{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=f_{\texttt{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\texttt{readout}}(\mathbf{H})$ for the entire graph. One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes. Indeed, these orderings do not hold any semantic meaning, and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max, or mean operators.
@@ 2.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework [1] which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
@@
As discussed above, we focus here on the functions that different GNN models employ to update node representations. Some GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [14, 15, 16]. For instance, each layer of the Graph Convolutional Network (GCN) [15] is defined as follows:
@@
However, the 1-layer perceptron is not a universal approximator of multiset functions [6], limiting the expressivity of the model. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [17, 18, 19, 7]. It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [6]. While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) [6], which is defined as follows:
@@
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned, which is not guaranteed. This has motivated a series of works that focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [20]. Other works have studied how to initialize the weights of the MLPs of the message-passing layers of GNNs. It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [21]. On the other hand, there exist settings where there is no need for complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [22] or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme [23].
@@ 2.3.1 Principle
Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community [12]. As mentioned above, this model relies on the Kolmogorov-Arnold representation theorem, which states that any multivariate function $f:\left[0,1\right]^{d}\rightarrow\mathbb{R}$ can be written as:
@@
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\texttt{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (6)
@@ 2.3.2 KAN Implementations and Applications
The original paper that introduced KANs uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture. Nonetheless, KANs allow for any function to serve as its base element. Thus, alternative implementations use different bases, such as radial-basis functions [24], which in particular alleviate some computational bottleneck of splines.
Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [25] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [26]. So far, most efforts have focused on time series data [27]. For instance, KANs have been evaluated in the satellite traffic forecasting task [28]. Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as the LSTM Network [29] and the Transformer [30]. The work closest to ours is the one reported in [31], where the authors propose FourierKAN-GCF. This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.

### v4

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message-passing) layers, within which the representation of each node is updated based on those of its neighbors. The most expressive message-passing GNNs can be obtained through the use of the sum aggregator and of MLPs for feature transformation, thanks to their universal approximation capabilities. However, the limitations of MLPs recently motivated the introduction of another family of universal approximators, called Kolmogorov-Arnold Networks (KANs) which rely on a different representation theorem. In this work, we compare the performance of KANs against that of MLPs on graph learning tasks. We implement three new KAN-based GNN layers, inspired respectively by the GCN, GAT and GIN layers. We evaluate two different implementations of KANs using two distinct base families of functions, namely B-splines and radial basis functions. We perform extensive experiments on node classification, link prediction, graph classification and graph regression datasets. Our results indicate that KANs are on-par with or better than MLPs on all tasks studied in this paper. We also show that the size and training speed of RBF-based KANs is only marginally higher than for MLPs, making them viable alternatives. Code available at https://github.com/RomanBresson/KAGNN.
@@ 1 Introduction
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations (Gilmer et al., 2017). Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs (Wu et al., 2020). Most GNNs consist of a series of message-passing layers. Within a message-passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
@@
A lot of recent work has focused on investigating the expressive power of GNNs (Li and Leskovec, 2022). There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot (Morris et al., 2019). Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs (Chen et al., 2019). This line of work gave insights into the limitations of different models (Xu et al., 2019; Nikolentzos et al., 2020), but also led to the development of more powerful architectures (Morris et al., 2019; Maron et al., 2019; Morris et al., 2020a).
@@
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities (Cybenko, 1989; Hornik et al., 1989). The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Recently, Kolmogorov-Arnold Networks (KANs) (Liu et al., 2025) have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957) which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs with learnable activations and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions and in settings where regularity is expected.
@@
In this paper, we present a thorough empirical comparison of the performance between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our intuition is that KANs could exploit the expected regularity underlying the message-passing paradigm to provide comparatively good performance on graph tasks. Our work is orthogonal to prior work that studies the expressive power of GNNs since we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the models’ performance. We implement two layers, one based on the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) and one based on the Graph Isomorphism Network (GIN) (Xu et al., 2019). We evaluate the different GNN models on several standard node classification, link prediction, graph classification, and graph regression datasets.
@@
The rest of this paper is organized as follows. Section 3, provides an overview of the tasks we address in this paper, as well as a description of message-passing GNNs and Kolmogorov-Arnold networks. In Section 4, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network), KAGCN (Kolmogorov-Arnold Graph Convolution Network) and KAGAT (Kolmogorov-Arnold Graph ATtention Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In Section 5, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks, and in Section 6 we evaluate the time efficiency of our architectures. Finally, Section 7 presents perspectives for future work and Section 8 concludes the paper.
2 Related Work
Kolmogorov-Arnold Networks have recently gained significant attention in the machine learning community (Somvanshi et al., 2024). Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification (Cheon, 2024), for predicting the pressure and flow rate of flexible electrohydrodynamic pumps (Peng et al., 2024), and for rumor detection (Jiang et al., 2024). So far, most efforts have focused on time series data (Xu et al., 2024a; Barašin et al., 2024). For instance, KANs have been evaluated in the satellite traffic forecasting task (Vaca-Rubio et al., 2024). Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as LSTM networks (Genet and Inzirillo, 2024a) and the Transformer (Genet and Inzirillo, 2024b). Recently, KAN-based convolutional neural networks have also been proposed which consist of kernels made of learnable non-linear functions that use splines (Bodner et al., 2024). These neural networks have been utilized for intrusion detection (Wang et al., 2025). KANs have also paved their way into celebrated architectures such as PointNet (Kashefi, 2024). Recent work has focused on improving different aspects of KANs such as making them more parameter-efficient (Ta et al., 2025), robust to overfitting (Altarabichi, 2024), and extending them to the complex domain (Wolff et al., 2025). More importantly, a framework was recently introduced which allows a seamless interplay between KANs and science (Liu et al., 2024). Specifically, multiplication was incorporated as a built-in modularity of KANs. Furthermore, the work strengthened the connection between KANs and scientific problems by highlighting the identification of key features, modular structures, and symbolic formulas.
Note that our study is not the only work in the intersection of KANs and graph learning algorithms (Kiamari et al., 2024; De Carlo et al., 2024; Zhang and Zhang, 2024; Li et al., 2024). However, while this paper provides an extensive evaluation across different settings and datasets, other works focus on a single task (e. g., node classification) or are limited to small-sized datasets. (Kiamari et al., 2024) propose two variants of the GCN architecture, where the feed-forward layer which updates node representations is replaced by KAN layers. The two variants differ from each other in how the new node representations are produced. The one transforms node representations before aggregation (and is thus similar to our KAGCN model), while the other after aggregation. However, the two variants are only evaluated on a single node classification dataset. (De Carlo et al., 2024) propose a neighborhood aggregation layer where spline-based activation functions are applied to the representations of the neighbors before aggregation, while another spline-based activation function is applied to the output of the aggregation. A final spline-based transformation then produces the final node representations. While the proposed model is evaluated in different tasks (i. e., node classification, graph classification and link prediction), all the considered datasets are small-sized. In another recent study, (Zhang and Zhang, 2024) also propose a variant of the GCN architecture, where a KAN is employed to transform node representations after aggregation. The proposed model is only evaluated on a single dataset where a graph is constructed from one-dimensional acoustic signals for axial flow pump fault diagnosis. Recent work has also investigated the ability of GNNs that employ KAN layers to solve problems in chemo- and bio-informatics such as the prediction of properties of molecules (Li et al., 2024) and the prediction of the binding affinity of small molecules to protein targets (Ahmed and Sifat, 2024), respectively. Such models were also applied to the problem of graph collaborative filtering (Xu et al., 2024b).
@@ 3.1 Graph Learning Tasks
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn how to map nodes to their class labels, i. e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=f_{\text{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\text{readout}}(\mathbf{H})$ for the entire graph. One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes. Indeed, these orderings do not hold any semantic meaning, and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max, or mean operators.
@@ 3.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework (Gilmer et al., 2017) which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
@@
As discussed above, we focus here on the functions that different GNN models employ to update node representations. Some GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features (Duvenaud et al., 2015; Kipf and Welling, 2017; Zhang et al., 2018). For instance, each layer of the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) is defined as follows:
@@
However, the 1-layer perceptron is not a universal approximator of multiset functions (Xu et al., 2019), limiting the expressivity of the model. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations (Balcilar et al., 2021; Dasoulas et al., 2021; Murphy et al., 2019; Nikolentzos et al., 2020). It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism (Xu et al., 2019). While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) (Xu et al., 2019), which is defined as follows:
@@
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned, which is not guaranteed. This has motivated a series of works that focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training (Guo et al., 2022). Other works have studied how to initialize the weights of the MLPs of the message-passing layers of GNNs. It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks (Han et al., 2023). On the other hand, there exist settings where there is no need for complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation (Wu et al., 2019) or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme (Gasteiger et al., 2019).
Finally, the Graph Attention Network (GAT)(Veličković et al., 2017) leverages the multi-head attention mechanism among nodes to give each neighbor an adapted importance in the aggregation scheme.
$\mathbf{h}^{(\ell)}_{v}=\sum_{u\in\mathcal{N}(v)\cup\{v\}}\alpha^{(\ell)}_{v,u}\mathbf{W}^{(\ell)}\mathbf{h}_{u}^{(\ell-1)}$ (4)
where $\mathbf{W}^{(\ell)}$ denotes a single linear layer and
$\alpha_{uv}^{(\ell)}=\frac{\exp\left(\texttt{LeakyReLU}\left(\left[\mathbf{W}^{(\ell)}h^{(\ell-1)}_{u}+\mathbf{W}^{(\ell)}h^{(\ell-1)}_{v}\right]\right)\right)}{\sum_{u\in\mathcal{N}(v)\cup\{v\}}\left(\exp\left(\texttt{LeakyReLU}\left(\left[\mathbf{W}^{(\ell)}h^{(\ell-1)}_{u}+\mathbf{W}^{(\ell)}h^{(\ell-1)}_{v}\right]\right)\right)\right)}$ (5)
are the neighbor-wise attention weights for a given head.
@@ 3.3.1 Principle
Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community (Liu et al., 2025). As mentioned above, this model relies on the Kolmogorov-Arnold representation theorem, which states that any multivariate function $f:\left[0,1\right]^{d}\rightarrow\mathbb{R}$ can be written as:
@@
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\text{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (8)
@@ 3.3.2 KAN Implementations and Applications
The original paper that introduced KANs uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture. Nonetheless, KANs allow for any function to serve as its base element. Thus, alternative implementations use different bases, such as radial-basis functions (Li, 2024), which in particular alleviate some computational bottleneck of splines.